Carbex, Inc., produces cutlery sets out of high-quality wood and steel. The company makes a Standard set and a Deluxe set and sells them to retail department stores throughout the country. The Standard set sells for $94, and the Deluxe set sells for $109. The variable expenses associated with each set are given below.

Standard

Deluxe

Variable production costs

$

32.00

$

47.00

Sales commissions (32% of sales price)

$

30.08

$

34.88

The company’s fixed expenses each month are:

Advertising

$

122,000

Depreciation

$

26,800

Administrative

$

71,500

Mary Parsons, the financial vice president, watches sales commissions carefully and has noted that they have risen steadily over the last year. For this reason, she was shocked to find that even though sales have increased, profits for the current month—May—are down substantially from April. Sales, in sets, for the last two months are given below:

Standard

Deluxe

Total

April

5,700

3,700

9,400

May

2,700

6,700

9,400

Required:

1-a. Prepare contribution format income statement for April.

1-b. Prepare contribution format income statement for May.

3-a. Compute the break-even point in dollar sales for April.

3-b. Would the break-even point in May be higher or lower than the break-even point in April?
Answer #1
Answer to 1-A :
Contribution Format Income statement for April
Particulars Amount ($) ROUGH WORK
Sales : (per unit) (No. of units)
Standard                        5,35,800 94 5700
Deluxe                        4,03,300 109 3700
Total Sales                       9,39,100
Less : Variable Costs :
Variable production costs :
Standard                        1,82,400 32 5700
Deluxe                        1,73,900 47 3700
Variable Selling expenses :
Standard                        1,71,456 30.08 5700
Deluxe                        1,29,056 34.88 3700
Total Variable Costs                       6,56,812
Contribution Margin (Sales - V.C)                       2,82,288
Less : Fixed Costs                        2,20,300
Net Income                          61,988
Answer to 1-B :
Contribution Format Income statement for May
Particulars Amount ($) ROUGH WORK
Sales : (per unit) (No. of units)
Standard                        2,53,800 94 2700
Deluxe                        7,30,300 109 6700
Total Sales                       9,84,100
Less : Variable Costs :
Variable production costs :
Standard                           86,400 32 2700
Deluxe                        3,14,900 47 6700
Variable Selling expenses :
Standard                           81,216 30.08 2700
Deluxe                        2,33,696 34.88 6700
Total Variable Costs                       7,16,212
Contribution Margin (Sales - V.C)                       2,67,888
Less : Fixed Costs                        2,20,300
Net Income                          47,588
Answer to 3-A :
Calculation of weighted average contribution margin ratio for April
Sale Ratio for both products :
Standard = 5700/9400 = 61%
Deluxe = 3700/9400 = 39%
Therefore, weighted average selling price = (94*61%)+(109*39%) = 100$
Therefore, weighted average variable expense = (62.08*61%)+(81.88*39%) = 70$
Hence Contribution per unit = 30$
Weighted average Contribution margin ratio = 30/100 = 30%
Fixed Costs = 220,300 $
Break-even point in $ sales for April = Fixed Cost / Weighted Average Contribution margin ratio = 220,300/30% = 7,34,333 $.
Answer to 3-B :
Calculation of weighted average contribution margin ratio for May
Sale Ratio for both products :
Standard = 2700/9400 = 29%
Deluxe = 6700/9400 = 71%
Therefore, weighted average selling price = (94*29%)+(109*71%) = 104.65$
Therefore, weighted average variable expense = (62.08*29%)+(81.88*71%) = 76.138$
Hence Contribution per unit = 28.512$
Weighted average Contribution margin ratio = 28.512/104.65 = 27%
Fixed Costs = 220,300 $
Break-even point in $ sales for May = Fixed Cost / Weighted Average Contribution margin ratio = 220,300/27% = 8,15,926 $.
Therefore, we can see that Break-even point in May is higher than break-even point in April (815,926>7,34,333)